What is the probability that the total of two dice will be greater than 8, given
that the first die is a 6?
1
Expert's answer
2020-09-28T19:19:39-0400

The first die is already 6, so to satisfy the condition that sum is greater than 8, the second die must be {3,4,5,6}. We have 4 events that satisfy the condition, and 6 possible events (getting any number from 1 to 6). So the probability is

"\displaystyle P(A) = \frac{4}{6} = \frac{2}{3}"

Answer: 2/3